# CP-Prompt: Composition-Based Cross-modal Prompting for Domain-Incremental Continual Learning

### v1

@@ Abstract.
The key challenge of cross-modal domain-incremental learning (DIL) is to enable the learning model to continuously learn from novel data with different feature distributions under the same task without forgetting old ones. However, existing top-performing methods still cause high forgetting rates, by lacking intra-domain knowledge extraction and inter-domain common prompting strategy. In this paper, we propose a simple yet effective framework, CP-Prompt, by training limited parameters to instruct a pre-trained model to learn new domains and avoid forgetting existing feature distributions. CP-Prompt captures intra-domain knowledge by compositionally inserting personalized prompts on multi-head self-attention layers and then learns the inter-domain knowledge with a common prompting strategy. CP-Prompt shows superiority compared with state-of-the-art baselines among three widely evaluated DIL tasks. The source code is available at https://github.com/dannis97500/CP_Prompt.
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; October 28 - November 1, 2024; Melbourne, Australia.isbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Object recognition
@@ 1. Introduction
Cross-modal models have garnered significant attention due to their capability to process and integrate diverse types of data. However, these models often encounter the challenge of different domains data feature distributions in practical applications. Domain Incremental Learning (DIL) (van de Ven et al., 2022) is a special incremental learning task, where the learning model is trained on a sequence of domains over time, with each domain or task presenting new and potentially information, e.g. distributional shift (Amodei et al., 2016). Under this setting, the tasks in each domain remain the same and the testing sample does not know which domain it belongs to. A vivid example is shown in Figure 1, where the learned model was firstly trained with quickdraw-style pictures, and then tested to classify the same category under different styles, such as infographics, painting, and clipart. The key success of DIL algorithm is to adapt and learn from sequential domains without forgetting the knowledge it has acquired from previous ones.
@@
  • We further designed Prefix-One, which can incorporate domain context information into the self-attention layer, therey guiding the transformer to fully utilize domain knowledge at different semantic levels for DIL process.
@@ 2. Related Work
Compared with traditional approaches, prompt learning-based DIL methods have shown prominent advantages in both model performance and efficiency. For example, DyTox (Douillard et al., 2022) applies the vision transformer in continual computer vision tasks, eliminating the class token and devising personalized tokens for each task. L2P (Wang et al., 2022c) employs a learnable key/prompt mechanism to select prompts added into image tokens based on the similarity between the keys and tokens. It introduces the concept of a prompt pool, based on a key-value mechanism to learn specific prompts within domains and make inferences by selecting appropriate prompts from the pool. Recent approaches such as S-liPrompts (Wang et al., 2022a) achieve remarkable results which independently learn a set of prompts for each domain using prompt tuning only, but still overlooking shared knowledge between domains and leaving significant room for improvement in intra-domain training methods. Not only limited ti DIL tasks, the class incremental learning (CIL) methods also employ prompt to optimize models. Dual-Prompt (Wang et al., 2022b) incorporates both general and expert prompts embedded in the pre-trained model, aimed at preserving personalized knowledge for retaining global shared knowledge and category distribution. HiDe-Prompt (Wang et al., 2023) expands the distance of category distribution using a contrastive loss penalty term, and identifies different tasks by optimizing the output layers.
@@ 4.3. Personalized Prompts
In addition to the complementary formation of common prompts, personalized prompts are embedded across the transformer’s attention layers in the form of parameters, capturing semantics at different granularities.
@@
The embeddings $\boldsymbol{x}_{p}^{s}$ then undergo projection transformation by the transformer layers of the pre-trained model. Therefore, inserting personalized prompts $\boldsymbol{P}_{\mathcal{P},img}^{s}\in\mathbb{R}^{L_{PI}\times D}$ into MSA layers can help to instruct the attention-capturing domain-individual semantics and knowledge, where $L_{PI}$ denotes the length of the personalized prompts.
Inspired by Prefix-Tuning (Li and Liang, 2021), we incorporate prompts into the prefix section of data features. To apply domain context into the MSA layer, we add soft prompts to the $Key$ matrix, and multiply it by the $Query$ matrix containing only the original data, then derive attention score. The attention score calculates the original data and the contextual information related to the domain style. The attention score is then multiplied by the $Value$ matrix that is the same as the $Key$. The obtained output incorporates the contextual knowledge within the current domain, and this attention extension is named as Prefix-One, as presented in Figure 3. It enables the pre-trained model to consider the particularity of personalized domain style and achieve the effect of adapting to domain tasks with less training cost.
@@ 4.4. Overall Objective for CP-Prompt
During the inference stage, we adopt a simple yet effective unsupervised clustering, $K$-Means, as domain selector to assign model extracted features with $K$ domain centroids $\mathbb{F}={\{\boldsymbol{m}_{f}^{j}\}}_{j=1}^{s}$ as feature pool. Given a new arriving sample $\boldsymbol{x}_{new}$, the domain selector selects the most relevant personalized prompts for inference by measuring the distance between $\boldsymbol{x}_{new}$ and $\mathbb{F}$. Following the multimodal pre-training setting (Wang et al., 2022a), the model inferences prediction $\boldsymbol{z}^{s^{\prime}}$ is derived by freezing prompts parameters and untuned pre-trained model:
@@
The overall training process is formally described in Algorithm 1.
0:  $\mathcal{T}$: Pre-trained image model; $\mathcal{G}$: Pre-trained language model; $\mathcal{D}_{s}={\{x^{s,i},y^{s,i}\}}_{i=1}^{N}$: Training data; $Y_{emb}=\{y_{emb}^{j}\}$: Text class embeddings.
1:  Initialization: Common Image Prompt Pool:$\boldsymbol{P}_{\mathcal{C}}$; Personalized Image Prompt Pool: $\boldsymbol{P}_{\mathcal{P},img}$; Personalized Language Prompt Pool:$\boldsymbol{P}_{\mathcal{P},tex}$; Domain centroids: $\mathbb{F}$
2:  for  $s=1,2,...S$  do
3:     Initialize common prompt $\boldsymbol{P}_{\mathcal{C}}^{s}=\boldsymbol{P}_{\mathcal{C}}^{s-1}$ for domain s
4:     Initialize personalized image prompt $\boldsymbol{P}_{\mathcal{P},img}^{s}$ for domain s
5:     Initialize personalized language prompt $\boldsymbol{P}_{\mathcal{P},tex}^{s}$ for domain s
6:     $\boldsymbol{x}_{p}^{s}$leftarrow Eq. (3)
7:     $\boldsymbol{h}_{p}^{(R)}$leftarrow Eq. (4)
8:     $\boldsymbol{t}_{h}^{s}$leftarrow Eq. (8) and Eq. (9)
9:     Calculate class embeddings $Y_{emb}$ by pre-trained model $\mathcal{G}$
10:     Compute the prediction probability by Eq. (11)
11:     Compute the Cross-Entropy loss by Eq. (12)
12:     Update $\theta_{P_{\mathcal{C}}^{s}};\theta_{P_{\mathcal{P},img}^{s}};\theta_{P_{\mathcal{P},img}^{s}}$
13:  end for
14:  Calculate training sample features by $\mathcal{T}(x^{s,i})$
15:  Calculate domain clustering centroids ${\{\boldsymbol{m}_{f}^{j}\}}_{j=1}^{s}$ by K-Means
16:  Save centroids in the list $\mathbb{F}$
Algorithm 1 The training algorithm of the CP-Prompt.
@@ 4.5. Model Analysis
In DIL scenarios, the parameters in the pre-training model are fixed. We argure that the attention score for the prompts should be further weighted, to learn the relationship between pre-training knowledge and domain-specific one, rather than simply adding parameters to fit new domain knowledge. As shown in Figure 5(b) of the original manuscript, we compare the proposed tuning strategy, prefix-one, with the prefix-tuning under the same environment setting among the three benchmarks. The result shows that CP-Prompt achieves optimal performance, with a 2% improvement.
@@ 5.1. Experiment Setup
Baselines.
Evaluation Metrics.
We employ widely used two standard metrics in DIL, average classification accuracy (AA) and forgetting rate (AF) as the evaluation metrics for comparing the CP-Prompt and other baselines. Formally, let $A_{i,n}$ represent the model evaluation score (i.e., classification accuracy in our experiment) on the $i$-th domain after training on the $n$-th domain. AA and AF are measured as follows:
(13) $AA=\frac{1}{n}\sum_{i=1}^{n}A_{i,n},$
(14) $AF=\frac{1}{n-1}\sum_{i=1}^{n-1}{(\frac{1}{n-i-1}\sum_{j=i+1}^{n}{(A_{i,j}-A_{i,i})})}.$
It should be noted that AA evaluates the learning model’s overall absolute performance, while AF measures the model’s ability to overcome catastrophic forgetting. A successful DIL model should strive for high AA while maximizing AF towards 0.
@@
Table 4. Ablation results of CP-Prompt on three datasets
5.2. Experimental Results
Main Results.
@@ 5.2. Experimental Results
We further explore contrasting different prompt design schemes between Dual-Prompt and the propsoed CP-Prompt. As introduced above, Dual-Prompt embeds the General Prompt within the attention layer, while CP-Prompt embeds the Common Prompt before the attention layer. As shown in Figure 7, we adopted the Dino visualization scheme (Caron et al., 2021) to map the attention variation when the model continues to learn on different domains. The attention weights of the first three layers of both approaches are displayed. It is observed that the attentions of CP-Prompt keep close to the object for identification alone side with domain shifting, while the Dual-Prompt gradually lose correct focus. It is evident that for cross-domain learning processes, embedding multi-domain common knowledge before attention better preserves key information in each domain. Particularly in info-graphs with multiple informational elements, the CP-Prompt approach can more accurately identify primary characteristic information.

### v2

@@ Abstract.
The key challenge of cross-modal domain-incremental learning (DIL) is to enable the learning model to continuously learn from novel data with different feature distributions under the same task without forgetting old ones. However, existing top-performing methods still cause high forgetting rates, by lacking intra-domain knowledge extraction and inter-domain common prompting strategy. In this paper, we propose a simple yet effective framework, CP-Prompt, by training limited parameters to instruct a pre-trained model to learn new domains and avoid forgetting existing feature distributions. CP-Prompt captures intra-domain knowledge by compositionally inserting personalized prompts into multi-head self-attention layers and then learns the inter-domain knowledge with a common prompting strategy. CP-Prompt shows superiority compared with state-of-the-art baselines among three widely evaluated DIL tasks. The source code is available at https://github.com/dannis97500/CP_Prompt.
@@
journalyear: 2024copyright: acmlicensedconference: Proceedings of the 32nd ACM International Conference on Multimedia; October 28-November 1, 2024; Melbourne, VIC, Australiabooktitle: Proceedings of the 32nd ACM International Conference on Multimedia (MM ’24), October 28-November 1, 2024, Melbourne, VIC, Australiadoi: 10.1145/3664647.3681481isbn: 979-8-4007-0686-8/24/10ccs: Computing methodologies Object recognition
@@ 1. Introduction
Cross-modal models have garnered significant attention due to their capability to process and integrate diverse types of data. However, these models often encounter the challenge of different domains data feature distributions in practical applications. Domain Incremental Learning (DIL) (van de Ven et al., 2022) is a special incremental learning task, where the learning model is trained on a sequence of domains over time, with each domain or task presenting new and potentially information, e.g. distributional shift (Amodei et al., 2016). Under this setting, the tasks in each domain remain the same and the testing sample does not know which domain it belongs to.
A vivid example is shown in Figure 1, where the learned model was firstly trained with quickdraw-style pictures, and then tested to classify the same category under different styles, such as infographics, painting, and clipart. The key success of DIL algorithm is to adapt and learn from sequential domains without forgetting the knowledge it has acquired from previous ones.
@@
  • We further designed Prefix-One, which can incorporate domain context information into the self-attention layer, thereby guiding the transformer to fully utilize domain knowledge at different semantic levels for DIL process.
@@ 2. Related Work
Compared with traditional approaches, prompt learning-based DIL methods have shown prominent advantages in both model performance and efficiency. For example, DyTox (Douillard et al., 2022) applies the vision transformer in continual computer vision tasks, eliminating the class token and devising personalized tokens for each task. L2P (Wang et al., 2022c) employs a learnable key/prompt mechanism to select prompts added into image tokens based on the similarity between the keys and tokens. It introduces the concept of a prompt pool, based on a key-value mechanism to learn specific prompts within domains and make inferences by selecting appropriate prompts from the pool. Recent approaches such as S-liPrompts (Wang et al., 2022a) achieve remarkable results which independently learn a set of prompts for each domain using prompt tuning only, but still overlooking shared knowledge between domains and leaving significant room for improvement in intra-domain training methods. Not only limited in DIL tasks, the class incremental learning (CIL) methods also employ prompt to optimize models. Dual-Prompt (Wang et al., 2022b) incorporates both general and expert prompts embedded in the pre-trained model, aimed at preserving personalized knowledge for retaining global shared knowledge and category distribution. HiDe-Prompt (Wang et al., 2023) expands the distance of category distribution using a contrastive loss penalty term, and identifies different tasks by optimizing the output layers.
@@ 4.3. Personalized Prompts
In addition to the complementary formation of common prompts, personalized prompts are embedded across the transformer’s attention layers in the form of parameters, capturing semantics at various levels of granularity.
@@
The embeddings $\boldsymbol{x}_{p}^{s}$ then undergo projection transformation by the transformer layers of the pre-trained model. Therefore, inserting personalized prompts $\boldsymbol{P}_{\mathcal{P},img}^{s}\in\mathbb{R}^{L_{PI}\times D}$ into MSA layers can help to instruct the attention-capturing domain-individual semantics and knowledge, where $L_{PI}$ denotes the length of the personalized prompts. Inspired by Prefix-Tuning (Li and Liang, 2021), we incorporate prompts into the prefix section of data features. To apply domain context into the MSA layer, we add soft prompts to the $Key$ matrix, and multiply it by the $Query$ matrix containing only the original data, then derive attention score. The attention score calculates the original data and the contextual information related to the domain style. The attention score is then multiplied by the $Value$ matrix that is the same as the $Key$. The obtained output incorporates the contextual knowledge within the current domain, and this attention extension is named as Prefix-One, as presented in Figure 3. It enables the pre-trained model to consider the particularity of personalized domain style and achieve the effect of adapting to domain tasks with less training cost.
@@ 4.4. Overall Objective for CP-Prompt
During the inference stage, we adopt a simple yet effective unsupervised clustering, $K$-Means, as domain selector to assign model extracted features with $K$ domain centroids $\mathbb{F}={\{\boldsymbol{m}_{f}^{j}\}}_{j=1}^{s}$ as feature pool. Given a new arriving sample $\boldsymbol{x}_{new}$, the domain selector selects the most relevant personalized prompts for inference by measuring the distance between $\boldsymbol{x}_{new}$ and $\mathbb{F}$. Following the multimodal pre-training setting (Wang et al., 2022a), the model’s inferences prediction $\boldsymbol{z}^{s^{\prime}}$ is derived by freezing prompts parameters and untrained pre-trained model:
@@ 4.5. Model Analysis
In DIL scenarios, the parameters in the pre-training model are fixed. We argue that the attention score for the prompts should be further weighted, to learn the relationship between pre-training knowledge and domain-specific one, rather than simply adding parameters to fit new domain knowledge. As shown in Figure 5(b) of the original manuscript, we compare the proposed tuning strategy, prefix-one, with the prefix-tuning under the same environment setting among the three benchmarks. The result shows that CP-Prompt achieves optimal performance, with a 2% improvement.
@@ 5.1. Experiment Setup
Baselines.
Evaluation Metrics.
We employ widely used two standard metrics in DIL(Wang et al., 2022a), average classification accuracy (AA) and forgetting rate (AF) as the evaluation metrics for comparing the CP-Prompt and other baselines.
@@
5.2. Experimental Results
Main Results.
Table 4. Ablation results of CP-Prompt on three datasets
@@ 5.2. Experimental Results
We further explore contrasting different prompt design schemes between Dual-Prompt and the proposed CP-Prompt. As introduced above, Dual-Prompt embeds the General Prompt within the attention layer, while CP-Prompt embeds the Common Prompt before the attention layer. As shown in Figure 7, we adopted the Dino visualization scheme (Caron et al., 2021) to map the attention variation when the model continues to learn on different domains. The attention weights of the first three layers of both approaches are displayed. It is observed that the attentions of CP-Prompt keep close to the object for identification alone side with domain shifting, while the Dual-Prompt gradually lose correct focus. It is evident that for cross-domain learning processes, embedding multi-domain common knowledge before attention better preserves key information in each domain. Particularly in info-graphs with multiple informational elements, the CP-Prompt approach can more accurately identify primary characteristic information.
@@ 6. Conclusion
Acknowledgements.
This work was supported by National Natural Science Foundation of China under Grant 62076035, China Postdoctoral Science Foundation under Grant 2022M711814, and by the Science and Technology Project of State Grid Hebei Information and Telecommunication Branch under Grant SGHEXT00SJJS2200026.